[Ncep.nhc.nco_contacts] TSB Morning Rounds - Thu Jul 07, 2016

TSB Admin nhc.tsbadmin at noaa.gov
Thu Jul 7 14:54:29 UTC 2016


Here is today's summary of NHC computer operations:
--- Craig Mattocks

----------

TAFB:

HSU:
1. ATCF: TAFB fixes did not import at 12 and 18 UTC on Wed, July 6. They  
were entered manually. ATCF dataflow expert Monica Bozeman is working on  
parallelizing the fix staging script so it processes multiple basins  
simultaneously, which should result in more timely fix imports.

2. ATCF: The primary disk partition on the secondary ATCF server (atcfsvr2)  
has reached 92% capacity. NCO SysAdmin Paul Obe is working on this problem.
Other items:
1. A kernel panic on the NESDIS primary FTP server,  
ftp.star.nesdis.noaa.gov, caused a 94-minute (11:21 to 12:55 UTC) outage of  
Himawari satellite data yesterday.

2. There was an emergency powerdown of the Luna and Tide WCOSS  
supercomputing clusters yesterday due to high temperatures (3 degrees above  
acceptable limits) at the Reston, VA facility caused by chiller issues.  
Service was restored 9 hours later.
NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160707/6fe4b3bc/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list